# ELT-Bench-Verified: Benchmark Quality Issues Underestimate AI Agent Capabilities

### 31 Mar 2026 v1

@@ 3. Methodology
Figure 5. Overview of the Auditor-Corrector framework. The Auditor proceeds in three phases: Phase 1 constructs a structured analysis environment for each failed task from the evaluation log; Phase 2 deploys an LLM agent per task that autonomously reverse-engineers the correct SQL for each unmatched column, verifying each derivation against the ground truth; Phase 3 involves manual validation and error categorization of all 660 reports. The Corrector applies category-specific corrections—evaluation script refinements and removal of columns with unreliable ground truth—to produce ELT-Bench-Verified.
@@ 3.1.2. Phase 2: Scaled Analysis via LLM Agent
3.1.3. Phase 3: Manual Verification and Categorization
A data engineer manually reviews all 660 analysis reports. For each report, the reviewer verifies that (i) the identified root cause is consistent with the presented evidence, (ii) the corrected SQL logic (if proposed) is semantically sound, and (iii) the match statistics are accurate.
Following validation, the reviewer assigns each column to an error category. The resulting categorization assigns each column $c\in\mathcal{U}$ to exactly one error category along two dimensions: attribution (agent-attributable vs. benchmark-attributable) and mitigability (whether the error can be corrected without modifying ground truth data). The primary categorization is performed by a single data engineer. To validate reproducibility, we conduct an inter-annotator agreement (IAA) study in which three annotators independently classify a uniformly sampled subset of 50 columns using the full 14-category taxonomy, given only the analysis reports and a detailed annotation protocol with category definitions, decision rules and worked examples. At the high level—which directly underpins the headline attribution statistics—Fleiss’ $\kappa=0.851$ (almost perfect agreement, 94.7%). At exact-category granularity across all 14 taxonomy classes, Fleiss’ $\kappa=0.755$ (substantial agreement, 79.3%), confirming that the fine-grained categorization is reproducible across annotators.
@@ 3.1.4. Validation of the Auditor
The structured prompt instructs the agent to produce, for each unmatched column, a dedicated analysis report and a Python debug script. Because the agent operates autonomously, its intermediate artifacts exhibit variation: debug scripts were sometimes written in SQL rather than Python (9 of 81 tasks), omitted entirely when the root cause was evident from direct inspection (5 tasks), or produced in multiple iterative versions as the agent refined its hypotheses (27 tasks). In 4 tasks where a single systemic root cause affected all columns in a table, the agent produced a combined analysis report rather than separate per-column reports. Despite these deviations, every unmatched column is covered: the 660 columns are addressed across 528 analysis reports, with multi-column reports accounting for the difference. Crucially, for 630 of 660 columns the agent’s corrected SQL is self-validated: the agent programmatically executes the corrected query against the source data and confirms a 100% row-level match with the ground truth before finalizing the report. For the remaining 30 columns, the agent explicitly reports that no SQL interpretation achieves a perfect match, flagging them as potential ground truth errors. A data engineer subsequently reviewed all 660 reports; no report required correction of its core diagnostic content.
@@ 3.2.2. Validation of the Corrector
Stratum B is the decisive stratum: these are the columns the patch reclassified from mismatch to match. We assess concordance between the consensus human labels and each script using McNemar’s test with Yates’ continuity correction throughout, which evaluates whether disagreements are symmetric or systematically biased in one direction. For Stratum B, the original script disagrees with human judgment on all 20 columns, always in the same direction—labeling columns as mismatches that humans consider equivalent—a bias that is highly statistically significant (McNemar $\chi^{2}=18.05$, $p<0.0001$). The patched script agrees with human judgment on all 20 Stratum B columns, with no discordant pairs. For Stratum A, both scripts achieve perfect concordance with human labels. For Stratum C, neither script agrees with human consensus on any of the 15 columns. The discordant pairs are nearly balanced ($b=7$, $c=8$), yielding McNemar $\chi^{2}=0.00$ ($p=1.0$) with continuity correction, indicating that neither script exhibits a systematic directional bias on these hard cases. The 0% agreement on Stratum C reflects the inherent difficulty of these cases: unlike the clear-cut representational equivalences corrected in Stratum B (e.g., boolean normalization, format differences), Stratum C columns involve subtler semantic judgments on which even human annotators do not fully agree. The fact that 7 of these 15 columns are labeled as equivalent by humans but as mismatches by both scripts suggests that our corrections are conservative by design—we address only the false positives for which strong consensus exists, and any residual cases would further strengthen the reported improvement. Over the stratified sample (which oversamples Stratum B to maximize statistical power on the decisive stratum), the patched script agrees with human judgment on 70% of sampled columns versus 30% for the original; note that per-stratum results are more informative than these unweighted aggregates, and its remaining disagreements show no directional bias (McNemar $\chi^{2}=0.00$, $p=1.0$), in contrast to the original script’s strongly asymmetric error pattern (McNemar $\chi^{2}=10.31$, $p=0.001$). These results confirm that the patch corrects genuine false positives validated by human judgment, without introducing new systematic errors. Table 1 presents the full per-stratum results.
Table 1. Stratified annotation study results (Fleiss’ $\kappa=0.664$, substantial). Concordance is measured against the majority-vote consensus label via McNemar’s test (1 d.f.). N/A = no discordant pairs; $^{***}p<0.001$; $^{**}p<0.01$. Overall figures are unweighted across strata; per-stratum results are more informative given deliberate oversampling of Stratum B.
Stratum Script Agree McNemar $\chi^{2}$
A ($n=15$) Original 15/15 (100%) N/A
Patched 15/15 (100%) N/A
B ($n=20$) Original 0/20 ( 0%) $18.05^{***}$
Patched 20/20 (100%) N/A
C ($n=15$) Original 0/15 ( 0%) 0.00(p=1.0)0.00\ (p{=}1.0)
Patched 0/15 ( 0%) 0.00(p=1.0)0.00\ (p{=}1.0)
Overall ($n=50$) Original 15/50 (30%) $10.31^{**}$
Patched 35/50 (70%) 0.00(p=1.0)0.00\ (p{=}1.0)
@@
Table 2. Validation study for the 30 columns categorized as having ground truth errors. Three engineers independently produced SQL and results are assessed via direct comparison. Low inter-annotator agreement indicates that the ground truth errors cannot be reliably corrected; low concordance with the auditor’s proposed values confirms no authoritative replacement can be established.
Comparison Exact-Match (%)
Inter-Annotator Agreement
Engineer A vs. Engineer B 70.0%
Engineer A vs. Engineer C 56.7%
Engineer B vs. Engineer C 46.7%
Average 57.8%
Auditor Concordance
Engineer A vs. Auditor 40.0%
Engineer B vs. Auditor 26.7%
Engineer C vs. Auditor 43.3%
Average 36.7%
Majority-Vote Concordance
Columns with consensus ($\geq$2/3) 24/30 (80.0%)
Majority vote vs. Auditor 10/24 (41.7%)
Table 2 presents the results. Pairwise inter-annotator agreement ranges from 46.7% to 70.0%, with an average of 57.8%. This level of disagreement among domain experts with full access to the source data indicates that no unambiguous correction exists for these columns. Concordance with the auditor’s proposed values is lower still, averaging 36.7% across engineers. Majority-vote consensus is available for only 24 of the 30 columns; on the remaining 6 columns, all three engineers produced entirely different results, further underscoring the fundamental ambiguity of these cases. Among the 24 columns with a majority, the consensus agrees with the auditor on only 41.7% of cases. The low inter-annotator agreement is the decisive criterion: it establishes that no single corrected ground truth can be asserted with confidence for these columns, rendering any revision arbitrary. Replacing the original erroneous values with the auditor’s proposals would therefore not constitute a correction but merely a substitution of one unreliable value for another. We consequently discard all 30 columns from the benchmark while retaining all remaining columns belonging to the same data model, amounting to a removal of just 30 out of 2,494 columns (1.2%) with no loss of data model coverage.
@@ 4.1.2. Benchmark-Attributable Errors
We define
benchmark-attributable errors as discrepancies arising from deficiencies in benchmark design rather than from agent reasoning. Specifically, these errors stem from ambiguous task specifications, limitations in the evaluation methodology, or incorrect ground truth values, leading to correct agent outputs being incorrectly classified as failures.
@@ 4.2. Error Distribution
Attribution Error Category Count %
Agent Flawed SQL Logic 221 33.5%
JOIN Type Errors 157 23.8%
Wrong Data Source 38 5.8%
Domain Knowledge Gaps 16 2.4%
Key Generation Errors 6 0.9%
NULL Handling Errors 4 0.6%
Subtotal (Agent) 442 67.0%
Benchmark Evaluation False Positives 156 23.6%
Ambiguous Data Model Descr. 32 4.8%
GT Calculation Errors 30 4.5%
Subtotal (Benchmark) 218 33.0%
Total 660 100%
@@
A total of 67 tasks (82.7% of failed tasks) contain at least one benchmark-attributable error. These flaws misdirect the community by penalizing correct behavior. We highlight two prominent categories:
@@
These are non-mitigable errors where the ground-truth data contains values that cannot be derived from any reasonable query over the available schema. For example, in the codebase_community database, a column defined as “set to 1 if the post was posted by the most influential user” has ground truth values that match only when “most influential” is operationalized as users with the specific reputation score of 995—ranking 306th out of all users—while users with reputations exceeding 87,000 are labeled as non-influential. This undocumented and counterintuitive rule cannot be deduced from the specification. As detailed in Appendix A.1, we tested five standard operationalizations of “most influential”—highest reputation, most posts, most badges, most profile views, and most upvotes—all of which are far more defensible, yet none reproduce the ground truth. Evaluating agents against such targets fundamentally breaks the utility of the benchmark.
@@ 5.2.1. End-to-End Re-evaluation
Benchmark SRDEL SRDT Models Passed
ELT-Bench (original) 96% 22.66% 46/203
ELT-Bench-Verified 96% 32.51% 66/203
@@ 5.2.3. Multi-Agent Validation
To demonstrate that the corrections are not overfit to a single agent, we evaluate a second agent configuration
—a Baseline agent implemented in LangGraph (LangChain, Inc., ) using the ReAct (Yao et al., 2023) agentic paradigm, paired with Claude Sonnet 4.5—on both the original ELT-Bench and ELT-Bench-Verified. Table 6 presents the results. Both agents show substantial improvement on ELT-Bench-Verified, confirming that the identified errors were genuine benchmark issues rather than artifacts of one agent’s behavior. Notably, both agents converge to the same SRDT (32.51%, 66/203 models) on the corrected benchmark despite differing on the original (22.66%, 46/203 models for SWE-Agent vs. 20.20%, 41/203 models for ReAct). However, the two agents arrive at the same 32.51% SRDT via different paths: ReAct achieves higher SRDEL (98% vs. 96%) and matches more columns overall (1,986 vs. 1,854), yet coincidentally passes the same number of data models (66/203). As shown in Appendix B, the 66 passing models are not identical across the two agents. The fact that both agents improve substantially on ELT-Bench-Verified—despite using different agentic strategies—provides evidence that the corrections address genuine benchmark defects rather than artifacts of a single agent’s behavior.

### 02 Apr 2026 v2

@@ 3. Methodology
Figure 5. Overview of the Auditor-Corrector framework. The Auditor proceeds in three phases: Phase 1 constructs a structured analysis environment for each failed task from the evaluation log; Phase 2 deploys an LLM agent per task that autonomously reverse-engineers the correct SQL for each unmatched column, verifying each derivation against the ground truth; Phase 3 involves manual categorization of all 660 reports into an error taxonomy. The Corrector then applies category-specific corrections—evaluation script refinements and removal of columns with unreliable ground truth—to produce ELT-Bench-Verified.
@@ 3.1.2. Phase 2: Scaled Analysis via LLM Agent
3.1.3. Phase 3: Categorization and Correction
A data engineer manually inspects all 660 analysis reports. For each report, the reviewer verifies that (i) the identified root cause is consistent with the presented evidence, (ii) the corrected SQL logic (if proposed) is semantically sound, and (iii) the match statistics are accurate.
Following inspection, the reviewer assigns each column to an error category. The resulting categorization assigns each column $c\in\mathcal{U}$ to exactly one error category along two dimensions: attribution (agent-attributable vs. benchmark-attributable) and mitigability (whether the error can be corrected without modifying ground truth data). The primary categorization is performed by a single data engineer. To ensure reproducibility, three independent annotators classified a uniformly sampled subset of 50 columns using the full 14-category taxonomy, given only the analysis reports and a detailed annotation protocol. The resulting agreement is strong at both granularity levels: for the high-level attribution (agent vs. benchmark), Fleiss’ $\kappa$ reaches 0.851 (almost perfect, 94.7% agreement), and for the exact 14-category classification, Fleiss’ $\kappa$ reaches 0.755 (substantial, 79.3% agreement), confirming that the categorization is reproducible across annotators.
@@ 3.1.4. Validation of the Auditor
The structured prompt instructs the agent to produce, for each unmatched column, a dedicated analysis report and a Python debug script. Because the agent operates autonomously, its intermediate artifacts exhibit natural variation: debug scripts were sometimes written in SQL rather than Python (9 of 81 tasks), or produced in multiple iterative versions as the agent refined its hypotheses (27 tasks). In 4 tasks where a single systemic root cause affected all columns in a table, the agent produced a combined analysis report rather than separate per-column reports. Despite these deviations, every unmatched column is covered: the 660 columns are addressed across 528 analysis reports, with multi-column reports accounting for the difference.
Crucially, for 630 of 660 columns the agent’s corrected SQL is self-validated: the agent programmatically executes the corrected query against the source data and confirms a 100% row-level match with the ground truth before finalizing the report. This execution-based verification ensures that each diagnosis is grounded in empirical evidence rather than reasoning alone. For the remaining 30 columns, the agent explicitly reports that no SQL interpretation achieves a perfect match, flagging them as potential ground truth errors—a finding later confirmed by the independent validation study described in Section 3.2.2. A data engineer subsequently reviewed all 660 reports, verifying the root cause identification, corrected SQL derivation, and supporting evidence in each case; no report required correction of its core diagnostic content.
@@ 3.2.2. Validation of the Corrector
Stratum B is the decisive stratum: these are the columns the patch reclassified from mismatch to match. We assess concordance between the consensus human labels and each script using McNemar’s test with Yates’ continuity correction throughout, which evaluates whether disagreements are symmetric or systematically biased in one direction. For Stratum B, the original script disagrees with human judgment on all 20 columns, always in the same direction—labeling columns as mismatches that humans consider equivalent—a bias that is highly statistically significant (McNemar $\chi^{2}=18.05$, $p<0.0001$). The patched script agrees with human judgment on all 20 Stratum B columns, with no discordant pairs. For Stratum A, both scripts achieve perfect concordance with human labels. For Stratum C, neither script agrees with human consensus on any of the 15 columns. The discordant pairs are nearly balanced ($b=7$, $c=8$), yielding McNemar $\chi^{2}=0.00$ ($p=1.0$) with continuity correction, indicating that neither script exhibits a systematic directional bias on these hard cases. The 0% agreement on Stratum C reflects the inherent difficulty of these cases: unlike the clear-cut representational equivalences corrected in Stratum B (e.g., boolean normalization, format differences), Stratum C columns involve subtler semantic judgments on which even human annotators do not fully agree. The fact that 7 of these 15 columns are labeled as equivalent by humans but as mismatches by both scripts suggests that our corrections are conservative by design—we address only the false positives for which strong consensus exists, and any residual cases would further strengthen the reported improvement. Over the stratified sample (which oversamples Stratum B to maximize statistical power on the decisive stratum), the patched script agrees with human judgment on 70% of sampled columns versus 30% for the original; note that per-stratum results are more informative than these unweighted aggregates, and its remaining disagreements show no directional bias (McNemar $\chi^{2}=0.00$, $p=1.0$), in contrast to the original script’s strongly asymmetric error pattern (McNemar $\chi^{2}=10.31$, $p=0.001$). These results confirm that the patch corrects genuine false positives validated by human judgment, without introducing new systematic errors.
@@
Pairwise inter-annotator agreement ranges from 46.7% to 70.0%, with an average of 57.8%. This level of disagreement among domain experts with full access to the source data indicates that no unambiguous correction exists for these columns. Concordance with the auditor’s proposed values is lower still, averaging 36.7% across engineers. Majority-vote consensus is available for only 24 of the 30 columns; on the remaining 6 columns, all three engineers produced entirely different results, further underscoring the fundamental ambiguity of these cases. Among the 24 columns with a majority, the consensus agrees with the auditor on only 41.7% of cases. The low inter-annotator agreement is the decisive criterion: it establishes that no single corrected ground truth can be asserted with confidence for these columns, rendering any revision arbitrary. Replacing the original erroneous values with the auditor’s proposals would therefore not constitute a correction but merely a substitution of one unreliable value for another. We consequently discard all 30 columns from the benchmark while retaining all remaining columns belonging to the same data model, amounting to a removal of just 30 out of 2,494 columns (1.2%) with no loss of data model coverage.
@@ 4. Audit Findings
Attribution Error Category Count %
Agent Flawed SQL Logic 221 33.5%
JOIN Type Errors 157 23.8%
Wrong Data Source 38 5.8%
Domain Knowledge Gaps 16 2.4%
Key Generation Errors 6 0.9%
NULL Handling Errors 4 0.6%
Subtotal (Agent) 442 67.0%
Benchmark Evaluation False Positives 156 23.6%
Ambiguous Data Model Descr. 32 4.8%
GT Calculation Errors 30 4.5%
Subtotal (Benchmark) 218 33.0%
Total 660 100%
@@ 4.1.2. Benchmark-Attributable Errors
We define benchmark-
attributable errors as discrepancies arising from deficiencies in benchmark design rather than from agent reasoning. Specifically, these errors stem from ambiguous task specifications, limitations in the evaluation methodology, or incorrect ground truth values, leading to correct agent outputs being incorrectly classified as failures.
@@ 4.2. Error Distribution
A total of 67 tasks (82.7% of failed tasks) contain at least one benchmark-attributable error. These flaws misdirect the community by penalizing correct behavior. We highlight the three benchmark-attributable categories with representative examples (full details in the Appendix):
@@
Ambiguous Data Model Descriptions (4.8%)
These arise when column specifications are underspecified, admitting multiple valid interpretations. For instance, the NUM_WON_2010 column in the european_football_2 database asks for “the number of matches the team won in the 2010 season.” In football, seasons span two calendar years, so “2010 season” could refer to the 2009/2010 season, the 2010/2011 season, or calendar year 2010. The agent selects 2009/2010—a defensible interpretation—yet the ground truth corresponds to 2010/2011, yielding only $\sim$40% agreement (Appendix A.3). We deliberately preserve such ambiguities in ELT-Bench-Verified, as resolving underspecified requirements is a capability agents should be evaluated on, not shielded from.
@@
These are non-mitigable errors where the ground-truth data contains values that cannot be derived from any reasonable query over the available schema. For example, in the codebase_community database, a column defined as “set to 1 if the post was posted by the most influential user” has ground truth values that match only when “most influential” is operationalized as users with the specific reputation score of 995—ranking 306th out of all users—while users with reputations exceeding 87,000 are labeled as non-influential. This undocumented and counterintuitive rule cannot be deduced from the specification. We tested five standard operationalizations—highest reputation, most posts, most badges, most profile views, and most upvotes—all far more defensible, yet none reproduce the ground truth (Appendix A.1). Evaluating agents against such targets fundamentally breaks the utility of the benchmark.
@@ 5.1. Corrections Applied
Benchmark SRDEL SRDT Models Passed
ELT-Bench (original) 96% 22.66% 46/203
ELT-Bench-Verified 96% 32.51% 66/203
@@ 5.2.3. Multi-Agent Validation
To demonstrate that the corrections are not overfit to a single agent, we evaluate a second agent
configuration—a Baseline agent implemented in LangGraph (LangChain, Inc., ) using the ReAct (Yao et al., 2023) agentic paradigm, paired with Claude Sonnet 4.5—on both the original ELT-Bench and ELT-Bench-Verified. Table 4 presents the results. Both agents show substantial improvement on ELT-Bench-Verified, confirming that the identified errors were genuine benchmark issues rather than artifacts of one agent’s behavior. Notably, both agents converge to the same SRDT (32.51%, 66/203 models) on the corrected benchmark despite differing on the original (22.66%, 46/203 models for SWE-Agent vs. 20.20%, 41/203 models for ReAct). However, the two agents arrive at the same 32.51% SRDT via different paths: ReAct achieves higher SRDEL (98% vs. 96%) and matches more columns overall (1,986 vs. 1,854), yet coincidentally passes the same number of data models (66/203). As shown in Appendix B, the 66 passing models are not identical across the two agents. The fact that both agents improve substantially on ELT-Bench-Verified—despite using different agentic strategies—provides evidence that the corrections address genuine benchmark defects rather than artifacts of a single agent’s behavior.
@@ A.2. Evaluation False Positive
A.3. Ambiguous Data Model Description
We illustrate ambiguous data model descriptions with the
NUM_WON_2010 column in the european_football_2 database, described as: “The number of matches the team won in the 2010 season, with NULL values replaced with 0.”
The term “2010 season” is inherently ambiguous in a football context, where seasons span two calendar years. The agent interprets it as the 2009/2010 season (August 2009–May 2010), a defensible reading. However, the ground truth values correspond to a different interpretation—either the 2010/2011 season or a calendar-year 2010 filter. Table 8 shows the resulting mismatches.
Table 8. Season interpretation ambiguity for NUM_WON_2010. The agent uses the 2009/2010 season; the ground truth implies a different interpretation. Only $\sim$40% of teams match under the agent’s reading.
Team GT Pred Diff
Ruch Chorzow 10 16 -6
Jagiellonia Bialystok 14 11 ++3
Lech Poznan 15 19 -4
P. Warszawa 10 9 ++1
Cracovia 6 9 -3
S.C. Olhanense 7 5 ++2
FC Pacos de Ferreira 8 8 0
SM Caen 4 0 ++4
The specification does not clarify whether “2010 season” refers to the season starting in 2010 (2010/2011), the season ending in 2010 (2009/2010), or matches played in the calendar year 2010. The agent’s choice of 2009/2010 is a reasonable default, yet it produces only $\sim$40% agreement with the ground truth. Testing the 2010/2011 season yields a 100% match, confirming that the mismatch stems entirely from an underspecified temporal reference rather than flawed SQL logic.